# DRIFT: Decompose, Retrieve, Illustrate, then Formalize Theorems

### v1

@@ 4.3 Implementation Details
In this study, we evaluate Drift as a lightweight prompting strategy, leveraging the in-context learning capabilities of instruction-following models such as DeepSeek-3.1 (DeepSeek-AI, 2024), GPT-4.1 (OpenAI, 2025), and Claude-Opus-4 (Anthropic, 2025) to demonstrate Drift’s broad applicability without introducing task-specific biases from finetuning. While we use the same model for both decomposition and formalization for consistency, these modules are independent and could be replaced with specialized finetuned models.
Decompose. We construct a few-shot prompt using five expertly-verified examples from the Putnam benchmark (Tsoukalas et al., 2024) to decompose the informal statements; full details in Section A.1.1 and Section A.3.1. Each decomposed sub-query consists of a natural language description of a concept and its formal representation retrieved from parametric knowledge, as detailed in Section 3.1. Through this, we instruct the LLMs to decompose problems.
Retrieve. The retriever model is a dense passage retriever666Implementation details in Section A.1.2 (DPR) finetuned on Mathlib data to map informal queries to their formal dependencies (Liu et al., 2025). We pre-compute embeddings for all formal declarations in the relevant libraries (Mathlib for ProofNet and MiniF2F-test; con-nf for ConNF). This training setup establishes ConNF as an OOD benchmark for the retrieval module. For the baseline, we retrieve the top-5 premises based on the entire informal statement.
@@
Formalize Theorems. As described in Section 3.4, the formalization prompt combines the original informal statement with the retrieved premises and illustrative theorems. Each premise is presented with its full name, formal declaration, and source code. Each illustrative theorem is included as a pair of its informal and formal statements. This demonstrates both the informal-to-formal alignment and the concrete application of the premises within a theorem instance. To evaluate pass@k, we generate 10 formalizations for each problem. The prompt context, including the retrieved premises and the illustrative theorems, remains consistent across all 10 runs.
@@ 4.4 Evaluation Metrics
We ablate the evaluation of Drift in two stages: a) the intrinsic performance of dependency retrieval plus the selection of illustrative theorems, and b) the extrinsic performance of autoformalization.
Intrinsic Dependency Retrieval. The effectiveness of the decomposition is measured by its impact on the dependency retrieval task as the quality of the decomposed sub-queries directly impacts the relevance of the retrieved premises. We measure the quality of the retrieved premises against the oracle* dependencies using Precision, Recall, and their harmonic mean, F1-score.
Formalization Correctness. For extrinsic evaluation, we use Typecheck (TC) and BEq+. Typecheck measures syntactic correctness, indicating the percentage of the generated statements that are valid and can pass the compiler’s type checker (Lu et al., 2024; Azerbayev et al., 2023; Liu et al., 2025). For semantic correctness, we use BEq+ (Poiroux et al., 2025), a symbolic metric that measures the logical equivalence between a predicted formal statement and the ground-truth reference by using deterministic proof tactics to bidirectionally prove that each statement can be transformed into the other. For each metric, we assess performance using pass@1 and pass@10, where pass@ indicates that at least one of independent generations was successful.
@@ 5.1 Dependency Retrieval
| DeepSeek-3.1 | 34.67 | 39.39 | 36.88 |
We evaluate the effectiveness of the Decompose and Retrieve modules by looking at their impact on intrinsic performance in dependency retrieval. As detailed in Table 1, we compare Drift against a no-decomposition baseline that uses the same dense retriever but with the original informal statements as queries. This provides a direct comparison to Drift, which retrieves a similar number of premises by taking the union of the top-1 results for each of the 5.21 to 6.42 sub-queries generated by the Decompose module. The results show that decomposition provides a substantial performance improvement in both precision and recall. Averaged across all decomposer models, Drift achieves an absolute improvement of 13.34, 2.08, and 7.74 points over the baseline F1 score on the ProofNet, MiniF2F-test, and ConNF benchmarks, respectively. Regarding the choice of decomposer, we observe that while Claude-Opus-4 achieves the highest F1 scores on ProofNet (27.68%) and MiniF2F-test (3.83%), the performance variation among the LLMs is marginal. Our findings indicate that frontier LLMs are largely interchangeable for this task, as the top-performing models have a maximum F1 score difference of only 2.05% on any benchmark.
@@ 5.2 Formalization
| Dataset | Formalizer | Retrieval | TC@1 | BEq+@1 | TC@10 | BEq+@10 |
@@
On autoformalization, Drift consistently outperforms both the zero-shot baseline and the strong retrieval-augmented baseline DPR (RAuto) on ProofNet and ConNF benchmarks, both on Typecheck and BEq+ with pass@1 and pass@10 (see Table 2). Specifically, on ProofNet, GPT-4.1 with Drift achieves a BEq+ pass@10 of 21.93%, a 2.74% improvement over DPR (RAuto). This trend holds across models in our main evaluation, demonstrating that our decomposition-driven approach provides more effective context for the formalization task. A key insight from our results is that while the frontier models we tested are all highly proficient and largely interchangeable as query decomposers, this parity does not extend to the final formalization step. For instance, DeepSeek-3.1 generally outperforms GPT-4.1 in the zero-shot setting across all benchmarks, suggesting a stronger parametric knowledge for direct formalization. However, this trend reverses when retrieval is introduced, most significantly on the OOD ConNF benchmark. On ConNF, all models achieve low zero-shot BEq+ scores (10%), confirming a severe knowledge gap. Retrieval substantially improves performance, where GPT-4.1 consistently outperforms DeepSeek-3.1 across all metrics with different retrieval strategies. Drift provides a particularly substantial improvement, increasing GPT-4.1’s BEq@10 score by 55.57%, even surpasses the Oracle* baseline by 3.43%. We hypothesize this is because the Oracle* provides only necessary premises based on the reference statement, while the illustrative theorems selected with Drift provide crucial demonstrations of premise usage.
On the in-distribution ProofNet benchmark, the results are more nuanced. GPT-4.1 surpasses DeepSeek-3.1 on BEq+@10 when using Drift. This pattern suggests that GPT-4.1 is more adept at in-context synthesis, integrating and reasoning over retrieved information to construct formal statements, especially in unfamiliar domains. The distinction suggests that the two stages of our pipeline rely on distinct LLM capabilities. Decomposition leverages natural language reasoning whereas formalization demands advanced formal reasoning.
As we discussed in Section 4, the MiniF2F-test benchmark presents a distinct profile with an average of only 0.43 library dependencies. This limits the potential for retrieval-based improvements, as evidenced by the small gap between the zero-shot and Oracle* performance, e.g. a pass@10 gap of only 2.24% for GPT-4.1 and 0.89% for DeepSeek-3.1. Instead, this low-dependency regime reveals the models’ high sensitivity to the provided context, which can act as a distractor rather than an aid.
@@ 5.3 Ablation study
In order to isolate and measure the contribution of each component of Drift, we conducted a systematic ablation study (Table 3). As expected, removing the illustrative theorems (w/o Illustrate) significantly decreased the BEq+ score on ProofNet and ConNF, which confirms that demonstrative examples of premise usage are crucial for the formalization correctness beyond just the definitions of the formal objects. Intriguingly, additionally removing the Decompose module (w/o Decompose) does not further degrade performance and even leads to a slight recovery on ConNF and ProofNet in BEq+ score. We hypothesize this is because the baseline DPR retrieves a thematically homogeneous (though less precise) set of premises via single query retrieval, which may be less distracting than the precise but more diverse set retrieved via decomposition. This reveals a crucial synergy: the illustrative theorems act as a scaffold that helps the model navigate the diverse information from the decomposer.
The complex interaction is further highlighted by the observations on MiniF2F-test, where removing theorems improved Typecheck rate but reduced BEq+ scores that measures the logical correctness of the generated formal statement. This further supports the hypothesis that for the simpler, low-dependency problems in MiniF2F-test, adding more context can act as a distractor. Removing external context improves syntactic validity but degrades logical correctness of the generated formal statements. This sensitivity strongly motivates the need for more dynamic and adaptive retrieval strategies. Future work on agentic frameworks could selectively retrieve information, judge its utility and iterate based on the compiler feedback.
@@ 6 Conclusion
In this work, we introduced Drift, a framework that improves autoformalization by tackling two distinct challenges: the underlying complexity of queries and the lack of contextual usage. Our decomposition-driven retrieval addresses the former by breaking down the informal statement into sub-queries and conducting point-to-point retrieval of its formal dependencies. Concurrently, our theorem selector resolves the latter by providing illustrative examples to guide the utilization of retrieved premises in theorem instances. This dual approach substantially improves formalization correctness on complex and out-of-distribution benchmarks, demonstrating its effectiveness as a broadly generalizable and model-agnostic strategy. On simpler, low-dependency MiniF2F-test dataset, our method performs on-par with related methods. Our findings suggest future work to focus on dynamic and adaptive retrieval strategies, as well as on agentic frameworks that iteratively refine attempts based on compiler feedback.
@@ A.1 Implementation Details
A.1.1 Few-Shot Example Generation for Decomposition
To construct a robust set of few-shot demonstrations for our decomposition module, we strategically selected five problems from the Putnam benchmark (Tsoukalas et al., 2024). These problems were chosen to ensure diversity in both their mathematical domain and the number of underlying premises required for their proofs.
We decomposed the informal statement of each selected problem into its atomic logical components using a zero-shot prompting strategy with Claude-Opus-4 in think mode, guided by a carefully engineered instruction set. To ensure the correctness and logical atomicity of these decompositions, each one was manually verified.
@@ A.1.2 Retriever Finetuning Details
We finetuned the BGE-M3 retriever model (Chen et al., 2024) on the Mathlib 4.7 dataset to specialize it for dependency retrieval in formal mathematics. The fine-tuning process was executed using the FlagEmbedding library (Chen et al., 2024) on a server equipped with four 32GB GPUs. The complete set of hyperparameters used for this process, including optimizer settings and loss configuration, is provided in Table 4.
@@ A.1.3 LLM Generation Parameters
For all generative tasks, sub-query generation (decomposition) and formal statement generation (formalization), we set the temperature to 0.7 for all models (GPT-4.1, DeepSeek-3.1, and Claude-Opus-4) to encourage diverse yet coherent outputs. To ensure reproducibility, single-attempt evaluations (Pass@1) used a fixed seed of 42. For multi-attempt evaluations (Pass@10), we generated ten distinct outputs by using a sequential range of seeds from 42 to 51.
@@ A.2.2 Dependency Retrieval Results of RAuto
This section presents the performance of DPR (RAuto) across both in-distribution (ProofNet, MiniF2F-test) and out-of-distribution (ConNF) benchmarks. The results, detailed in Table 5, highlight a crucial trade-off between specialization and generalization that motivates our proposed approach. When comparing the baselines, our baseline retriever substantially outperforms DPR (RAuto) on the ConNF benchmark but underperforms on ProofNet and MiniF2F-test. We hypothesize that this discrepancy arises because DPR (RAuto) may be overfitted to Mathlib-specific content.
| Dataset | Precision | Recall | F1 |
@@ A.2.3 Autoformalization Results of Claude
Table 6 summarizes the pass@1 results of Claude-Opus-4, reporting both the Typecheck rate (TC@1) and the equivalence rate (BEq+@1) across the ProofNet, MiniF2F-test, and ConNF benchmarks. These results provide context for the performance Claude-Opus-4 discussed in the main paper.
| Dataset | Retrieval | TC@1 | BEq+@1 |
@@ A.2.4 The Role of Illustrative Theorems as a Scaffold
As presented in Table 3, removing the query decomposer (w/o Decompose: reverting to the baseline DPR) does not degrade performance further. In fact, on ConNF and ProofNet, it led to a slightly recovery in the BEq+ score compared to the “w/o Illustrate” setting. We hypothesize this is due to the nature of retrieval noise. The baseline DPR, using a single query, retrieves a thematically clustered set of premises. While its precision is lower, its noise is homogeneous and may be less distracting to the LLM. Our decomposition method retrieves more diverse set of premises. While this captures more correct dependencies (higher recall and precision), the accompanying noise is also more varied. This reveals a crucial synergy: the theorem selector acts as a contextual scaffold, helping the model navigate the diverse information retrieved by the decomposer. Without this guidance, the varied noise can outweigh the benefit of improved retrieval.
@@ A.2.5 Scaling Performance with Sampling
Across all experiments in Table 2, we observe a consistent and significant gap between pass@1 and pass@10 results. For instance, performance on ProofNet improved by an average of 27.20% across all settings and formalizer models. This large uplift underscores the potential for enhancing performance through sampling-based methods at test time. This suggests that performance could be further scaled by integrating our method into an agentic framework equipped with a verifier (e.g Typecheck correctness).
Notably, DeepSeek-3.1’s performance of pass@10 saturates more quickly than Claude-Opus-4’s on the ProofNet benchmark. Its zero-shot Pass@10 score is only 0.27% lower than its retrieval-augmented Drift score. This suggests that with sufficient sampling, the model can sometimes recover the necessary knowledge parametrically. We anticipate a similar, albeit slower, trend for the larger Claude-Opus-4 and GPT-4.1 models if the number of attempts were increased further.

### v3

@@ 4.3 Implementation Details
In this study, we evaluate Drift as a lightweight prompting strategy, leveraging the in-context learning capabilities of instruction-following models such as DeepSeek-V3.1 (DeepSeek-AI, 2024), GPT-4.1 (OpenAI, 2025), and Claude-Opus-4 (Anthropic, 2025a) to demonstrate Drift’s broad applicability without introducing task-specific biases from finetuning. While we use the same model for both decomposition and formalization for consistency, these modules are independent and could be replaced with specialized finetuned models.
Decompose. We construct a few-shot prompt using five expertly-verified examples from the Putnam benchmark (Tsoukalas et al., 2024) to instruct LLMs to decompose the informal statements; full details in Section A.1.1 and Section A.3.1. Each decomposed sub-query consists of a natural language description of a concept and its formal representation predicted by the decomposer with parametric knowledge, as detailed in Section 3.1.
Retrieve. The retriever model is a dense passage retriever666Implementation details in Section A.1.2. (DPR) finetuned on Mathlib data to map informal queries to their formal dependencies (Liu et al., 2025). We pre-compute embeddings for all formal declarations in the relevant libraries (Mathlib for ProofNet and MiniF2F-test; con-nf for ConNF). This training setup establishes ConNF as an OOD benchmark for the retrieval module. For the DPR baseline, we retrieve the top-5 premises based on the entire informal statement.
@@
Formalize Theorems. As described in Section 3.4, the formalization prompt combines the original informal statement with the retrieved premises and illustrative theorems. Each premise is presented with its full name, formal declaration, and source code. Each illustrative theorem is included as a pair of its informal and formal statements. This demonstrates both the informal-to-formal alignment and the concrete application of the premises within a theorem instance. To evaluate pass@k, we generate 10 formalizations for each problem.
@@ 4.4 Evaluation Metrics
We conduct the evaluation of Drift in two stages: a) the intrinsic performance of dependency retrieval and the selection of illustrative theorems, and b) the extrinsic performance of autoformalization.
Dependency Retrieval. The effectiveness of the decomposition is measured by its impact on the dependency retrieval task as the quality of the decomposed sub-queries directly impacts the relevance of the retrieved premises. We measure the quality of the retrieved premises against the oracle* dependencies using Precision, Recall, and their harmonic mean, F1-score.
Formalization Correctness. For formalization, we use Typecheck (TC) and BEq+. Typecheck measures syntactic correctness, indicating the percentage of the generated statements that are valid and can pass the compiler’s type checker (Lu et al., 2024; Azerbayev et al., 2023; Liu et al., 2025). For semantic correctness, we use BEq+ (Poiroux et al., 2025), a symbolic metric that measures the logical equivalence between a predicted formal statement and the ground-truth reference by using deterministic proof tactics to bidirectionally prove that each statement can be transformed into the other. For each metric, we assess performance using pass@1 and pass@10, where pass@ indicates that at least one of independent generations was successful.
@@ 5.1 Dependency Retrieval
| DeepSeek-V3.1 | 34.67 | 39.39 | 36.88 |
We evaluate the effectiveness of the Decompose and Retrieve modules by looking at their impact on intrinsic performance in dependency retrieval. As detailed in Table 1, we compare Drift against a monolithic query baseline that uses the same dense retriever but with the original informal statements as queries. This provides a direct comparison to Drift, which retrieves a similar number of premises by taking the union of the top-1 results for each of the 5.21 to 6.42 sub-queries generated by the Decompose module.777The full statistics of the decomposed sub-queries are available at Section A.2.7. The results show that decomposition provides a substantial performance improvement in both precision and recall. Averaged across all decomposer models, Drift achieves an absolute improvement of 13.34, 2.08, and 7.74 points over the baseline F1 score on the ProofNet, MiniF2F-test, and ConNF benchmarks, respectively. Regarding the choice of decomposer model, we observe that while Claude-Opus-4 achieves the highest F1 scores on ProofNet (27.68%) and MiniF2F-test (3.83%), the performance variation among the LLMs is marginal. Our findings indicate that frontier LLMs are largely interchangeable for this task as the top-performing models have a maximum F1 score difference of only 2.05% on any benchmark.
@@ 5.2 Formalization
| Benchmark | Formalizer | Retrieval | TC@1 | BEq+@1 | TC@10 | BEq+@10 |
@@
For extrinsic performance on autoformalization, Drift consistently outperforms both the zero-shot baseline and the strong retrieval-augmented baseline DPR (RAuto) on ProofNet and ConNF benchmarks, both on Typecheck and BEq+ with pass@1 and pass@10 (see Table 2). Specifically, on ProofNet, GPT-4.1 with Drift achieves a BEq+ pass@10 of 21.93%, a 2.74% improvement over DPR (RAuto). This trend holds across models in our main evaluation, demonstrating that our decomposition-driven approach provides more effective context for the formalization task.
A key insight from our results is that while the frontier models we tested are all highly proficient and largely interchangeable as query decomposers, this parity does not extend to the final formalization step. For instance, DeepSeek-V3.1 generally outperforms GPT-4.1 in the zero-shot setting across all benchmarks, suggesting a stronger parametric knowledge for direct formalization. However, this trend reverses when retrieval is introduced, most significantly on the OOD ConNF benchmark. On ConNF, all models achieve low zero-shot BEq+ scores (10%), confirming a severe knowledge gap. Retrieval substantially improves performance, where GPT-4.1 consistently outperforms DeepSeek-V3.1 across all metrics with different retrieval strategies. Drift provides a particularly significant improvement, increasing GPT-4.1’s BEq+@10 score by 55.57% and even surpassing the oracle* baseline by 3.43%. We hypothesize this is because the oracle* provides only necessary premises based on the reference statement, while the illustrative theorems selected with Drift provide crucial demonstrations of premise usage.
On the in-distribution ProofNet benchmark, the results are more nuanced. GPT-4.1 surpasses DeepSeek-V3.1 on BEq+@10 when using Drift. This pattern suggests that GPT-4.1 is more adept at in-context synthesis, integrating and reasoning over retrieved information to construct formal statements, especially in unfamiliar domains. In contrast to the interchangeability we observed among models as decomposers, the formalization stage reveals clear performance differences. The distinction suggests that the two stages of our pipeline rely on distinct LLM capabilities. Decomposition leverages natural language reasoning, whereas formalization demands advanced formal reasoning.
As we discussed in Section 4, the MiniF2F-test benchmark presents a distinct profile with an average of only 0.43 library dependencies. This limits the potential for retrieval-based improvements, evidenced by the small gap between the zero-shot and oracle* performance, (e.g., a pass@10 gap of 2.24% for GPT-4.1 and 0.89% for DeepSeek-V3.1). Instead, this low-dependency regime reveals the models’ high sensitivity to the provided context, which can act as a distractor rather than an aid.
@@ 5.3 Ablation study
In order to isolate and measure the contribution of each component of Drift, we conducted a systematic ablation study (Table 3). As expected, removing the illustrative theorems (w/o Illustrate) decreased the BEq+ score on ProofNet and ConNF, which confirms that demonstrations of premise usage are crucial for the formalization correctness beyond just the definitions of the formal objects. Intriguingly, additionally removing the Decompose module (w/o Decompose) does not further degrade performance and even leads to a slight recovery on ConNF and ProofNet in BEq+ score. We hypothesize this is because the baseline DPR retrieves a thematically homogeneous (though less precise) set of premises via single query retrieval, which may be less distracting than the precise but more diverse set retrieved via decomposition. This reveals a crucial synergy: the illustrative theorems act as a scaffold that helps the model navigate the diverse information from the decomposer.
The complex interaction creates a trade-off on the MiniF2F-test benchmark: removing theorems improves syntactic correctness (Typecheck) while degrading logical correctness (BEq+). This further supports the hypothesis that for the simpler, low-dependency problems in MiniF2F-test, adding more context can act as a distractor. Removing external context improves syntactic validity but degrades logical correctness of the generated formal statements. This sensitivity strongly motivates the need for more dynamic and adaptive retrieval strategies. Future work on agentic frameworks could selectively retrieve information, judge its utility, and iterate based on the compiler feedback.
@@ 6 Conclusion
In this work, we introduced Drift, a framework that improves autoformalization by tackling two distinct challenges: the underlying complexity of queries and the lack of contextual usage. Our decomposition-driven retrieval addresses the former by breaking down the informal statement into sub-queries and conducting point-to-point retrieval of its formal dependencies. Concurrently, the Illustrate module resolves the latter by providing illustrative examples to guide the utilization of retrieved premises in theorem instances. This dual approach substantially improves formalization correctness on both complex in-distribution (ProofNet) and out-of-distribution (ConNF) benchmarks, demonstrating its effectiveness as a broadly generalizable and model-agnostic strategy. On a simpler, low-dependency MiniF2F-test benchmark, our method performs comparably to related methods. Our findings suggest future work to focus on dynamic and adaptive retrieval strategies, as well as on agentic frameworks that iteratively refine attempts based on compiler feedback.
@@ A.1 Implementation Details
A.1.1 Few-Shot Examples for Decomposition
To construct a robust set of few-shot demonstrations for our Decompose module, we strategically selected five problems from the Putnam benchmark (Tsoukalas et al., 2024). These problems were chosen to ensure diversity in both their mathematical domain and the number of underlying premises required for their formal statements.
We decomposed the informal statement of each selected problem into its atomic, logical sub-components using Claude-Opus-4 with extended thinking enabled (Anthropic, 2025b). The decomposition followed a zero-shot prompting strategy guided by a carefully engineered instruction set. To ensure correctness and logical atomicity, each decomposition was manually verified.
@@ A.1.2 Retriever Finetuning Details
We finetuned the BGE-M3 retriever model (Chen et al., 2024) on the Mathlib 4.7 dataset to specialize it for dependency retrieval in formal mathematics. The finetuning process was executed using the FlagEmbedding library (Chen et al., 2024) on a server equipped with four 32GB GPUs. The complete set of hyperparameters used for this process, including optimizer settings and loss configuration, is provided in Table 4.
@@ A.1.3 LLM Generation Parameters
For all generative tasks, sub-query generation (decomposition) and formal statement generation (formalization), we set the temperature to 0.7 for all models (GPT-4.1, DeepSeek-V3.1, and Claude-Opus-4) to encourage diverse yet coherent outputs. To ensure reproducibility, single-attempt evaluations (pass@1) used a fixed seed of 42. For multi-attempt evaluations (pass@10), we generated ten distinct outputs by using a sequential range of seeds from 42 to 51.
@@ A.2.2 Dependency Retrieval Results of RAuto
This section presents the performance of DPR (RAuto) across both in-distribution (ProofNet, MiniF2F-test) and out-of-distribution (ConNF) benchmarks. The results, detailed in Table 5, highlight a crucial trade-off between specialization and generalization that motivates our proposed approach. When comparing DPR baselines, our retriever without decomposition substantially outperforms DPR (RAuto) on the ConNF benchmark but underperforms on ProofNet and MiniF2F-test. We hypothesize that this discrepancy arises because DPR (RAuto) may be overfitted to Mathlib-specific content.
| Benchmark | Precision | Recall | F1 |
@@ A.2.3 Autoformalization Results of Claude
In the main paper, we evaluated Claude-Opus-4’s effectiveness as a decomposer for query decomposition. Here, we provide Claude-Opus-4’s formalization results for completeness. Table 6 reports pass@1 performance (TC@1 and BEq+@1) across all benchmarks. Note that pass@10 results for Claude-Opus-4 are not available due to computational costs.
| Benchmark | Retrieval | TC@1 | BEq+@1 |
@@ A.2.4 The Role of Illustrative Theorems as a Scaffold
As presented in Table 3, removing the Decompose module (w/o Decompose: reverting to the baseline DPR) does not degrade performance further. In fact, on ConNF and ProofNet, it leads to a slight recovery in the BEq+ score compared to the “w/o Illustrate” setting. We hypothesize this is due to the nature of retrieval noise. The baseline DPR, using a single query, retrieves a thematically clustered set of premises. While its precision is lower, its noise is homogeneous and may be less distracting to the LLM. Our decomposition method retrieves a more diverse set of premises. While this captures more correct dependencies (higher recall and precision), the accompanying noise is also more varied. This reveals a crucial synergy: the selected theorems in the Illustrate module act as a contextual scaffold, helping the model navigate the diverse information retrieved by the decomposer. Without this guidance, the varied noise can outweigh the benefit of improved retrieval.
@@ A.2.5 Scaling Performance with Sampling
Across all experiments in Table 2, we observe a consistent and significant gap between pass@1 and pass@10 results. For instance, performance on ProofNet improved by an average of 27.20% across all settings and formalizer models. This large uplift underscores the potential for enhancing performance through sampling-based methods at test time. This suggests that performance could be further scaled by integrating our method into an agentic framework equipped with a verifier (e.g., Typecheck correctness).
Notably, DeepSeek-V3.1’s performance of pass@10 saturates more quickly than Claude-Opus-4’s on the ProofNet benchmark. Its zero-shot pass@10 score is only 0.27% lower than the Drift score. This suggests that with sufficient sampling, the model can sometimes recover the necessary knowledge parametrically. We anticipate a similar, albeit slower, trend for the larger Claude-Opus-4 and GPT-4.1 models if the number of attempts were increased further.